# GeneralizIT: A Python Solution for Generalizability Theory Computations thanks: Citation:

### v1

@@ 1 Statement of Need
G_String_MVb Free from GitHub Raw data Yes Yes Yes
SAS VARCOMP Included in SAS Raw data Yesc Yes No
SPSS VARCOMP Included in SPSS Raw data Yesd Yes No
@@ 3.1.1 Calculating Sum of Squares and Mean Squares for Variance Components
$MS(\alpha)=\frac{SS(\alpha)}{Do(\alpha)}$
@@ 4 Conclusion
GeneralizIT is a Python package that simplifies the application of Generalizability Theory in research and practice. By automating the calculation of variance components, generalizability and dependability coefficients, and conducting D-Studies, GeneralizIT empowers researchers to conduct robust reliability analyses with minimal coding effort in an analysis environment with which they are familiar. Currently, no Python package exists for conducting G-Theory analyses, creating a significant barrier to its widespread adoption. This work addresses this critical need by providing a user-friendly platform for balanced crossed and nested designs. Other G-Theory software solutions allow for missing data or more complex unbalanced designs however they are more user intensive. We reserve these features for future work. Presently, GeneralizIT offers an accessible, computationally efficient solution for improving measurement reliability in diverse research environments. By democratizing the use of G-Theory, GeneralizIT provides a valuable resource for researchers seeking to enhance the validity and reliability of their measurement instruments and make sound statistically informed decisions.

### v2

@@ 1 Statement of Need
G_String_MVb Free from GitHub Raw data Noc Yes Yes
SAS VARCOMP Included in SAS Raw data Yesd Yes No
SPSS VARCOMP Included in SPSS Raw data Yese Yes No
@@
    Mean substitutions if missing data detected prior to analysis
  • d
@@ 3.1.1 Calculating Sum of Squares and Mean Squares for Variance Components
$MS(\alpha)=\frac{SS(\alpha)}{Df(\alpha)}$
@@ 4 Conclusion
GeneralizIT is a Python package that simplifies the application of Generalizability Theory in research and practice. By automating the calculation of variance components, generalizability and dependability coefficients, and conducting D-Studies, GeneralizIT empowers researchers to conduct robust reliability analyses with minimal coding effort in an analysis environment with which they are familiar. Currently, no Python package exists for conducting G-Theory analyses, creating a significant barrier to its widespread adoption. This work addresses this critical need by providing a user-friendly platform for balanced crossed and nested designs. While other G-Theory software solutions allow for analyses that take into account missing data and/or work with unbalanced designs, they are more user intensive and/or proprietary. We reserve these features for future work. Presently, GeneralizIT offers an accessible, computationally efficient solution for improving measurement reliability in diverse research environments. By democratizing the use of G-Theory, GeneralizIT provides a valuable resource for researchers seeking to enhance the validity and reliability of their measurement instruments and make sound statistically informed decisions.
5 Acknowledgments
The authors would like to acknowledge Professor Theresa Kline for her support and insights into the methodology of Generalizability theory as they were invaluable to this work.